# New Convergence Analysis of GMRES with

### v1

@@ 9 Conclusions
We present a very general convergence bound for preconditioned GMRES when any inner product is used, and when it is deflated. We consider bounds for several generic cases for the deflation space. These bounds inspired us to produce an effective deflation space, namely, the eigenvectors of the generalized eigenvalue problem $\mathbf{N}\mathbf{z}=\lambda\mathbf{M}\mathbf{z}$, where $\mathbf{M}$ is the Hermitian part of $\mathbf{A}$, which is assumed to be positive definite, and $\mathbf{N}$ is the skew-Hermitian part of $\mathbf{A}$. Only eigenvectors corresponding to eigenvalues with muduli above a given threshold are selected. Numerical experiments illustrate the potential for these ideas. Deflating indeed reduces the number of iterations, and so do the preconditioners combined with the deflation. On the other hand, while our theory applies to any inner product, our experiments do not show an improvement with our choices of the weights.

### v2

@@ 8 Conclusions
We present a very general convergence bound for preconditioned GMRES when any inner product is used, and when it is deflated. We consider bounds for several generic cases for the deflation space. These bounds inspired us to produce an effective deflation space, namely, the eigenvectors of the generalized eigenvalue problem $\mathbf{N}\mathbf{z}=\lambda\mathbf{M}\mathbf{z}$, where $\mathbf{M}$ is the Hermitian part of $\mathbf{A}$, which is assumed to be positive definite, and $\mathbf{N}$ is the skew-Hermitian part of $\mathbf{A}$. Only eigenvectors corresponding to eigenvalues with moduli above a given threshold are selected. Numerical experiments illustrate the potential for these ideas. In some cases, deflating is a key to convergence when GMRES without deflation fails to converge. In other cases, deflating indeed reduces the number of iterations, and so do the preconditioners combined with the deflation. On the other hand, while our theory applies to GMRES in the preconditioner inner product, our experiments show that the choice of inner product is not as important as the deflation space.
Acknowledgements
We thank the referees for making useful comments and suggestions, which helped improved our presentation.